# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ D.1 Identifiable transformation of multinary-product tree
$\mathbb{E}_{\nu,W}(\gamma_{\boldsymbol{\ell}}):=\sum_{\boldsymbol{\ell}_{W}\in\prod_{t\in W}[\phi_{t}]}\gamma_{(\boldsymbol{\ell}_{W^{c}},\boldsymbol{\ell}_{W})}\nu_{W}^{\text{ind}}\{I_{\boldsymbol{\ell}_{W}}\}.$ (D.1)
@@
$f_{\nu,f_{S}}(\mathbf{x}_{S})=\sum_{\boldsymbol{\ell}}\tilde{\gamma}_{\boldsymbol{\ell}}\mathbb{I}(\mathbf{x}_{S}\in I_{\boldsymbol{\ell}}).$ (D.2)
D.2 Approximation error between the empirically and populationally identifiable multinary-product trees
We first investigate the approximation error between a given empirically identifiable multinary-product tree and its populationally identifiable version obtained by the formula (D.2), whose result is given in the following theorem. For $S\subseteq[p]$, let $\mathbb{P}_{S}$ be the distribution of $\mathbf{X}_{S}$ and $\mathbb{P}_{S}^{\text{ind}}=\prod_{j\in S}\mathbb{P}_{j}$, where $\mathbb{P}_{j}$ is the probability distribution of $\mathbf{X}_{j}$.
Theorem D.1.
@@ D.2 Approximation error between the empirically and populationally identifiable multinary-product trees
$=\sum_{\ell_{j}\in[\phi_{j}]}\gamma_{(\boldsymbol{\ell}_{j^{c}},\ell_{j})}\bigg{(}\mathbb{P}_{j}\{I_{j,\ell_{j}}\}-{1\over n}\sum_{i=1}^{n}\mathbb{I}(X_{i,j}\in I_{j,\ell_{j}})\bigg{)}.$ (D.3)
In the following, we will show that (D.3) is small.
@@
$\mathbb{P}_{\mathbf{X}}^{n}\bigg{\{}\sup_{q\in\mathcal{Q}_{K,\xi}}\bigg{|}{1\over n}\sum_{i=1}^{n}q(X_{i,j})-\mathbb{E}_{j}[q(X_{j})]\bigg{|}>(24\xi+1)\sqrt{K\log n\over n}\bigg{\}}\leq 2\exp\bigg{(}-{K\log n\over 4\xi^{2}}\bigg{)},$ (D.4)
where $\mathcal{Q}_{K,\xi}$ is the set of all $q_{\mathcal{P},\boldsymbol{\gamma}}$ with $\phi_{j}\leq K$ and $\boldsymbol{\gamma}\in[-\xi,\xi]^{\phi_{j}},$ and $\mathbb{E}_{j}$ denotes the expectation under $\mathbb{P}_{j}$.
@@
$\bigg{|}\sup_{q\in\mathcal{Q}_{K,\xi}}\triangle_{\mathbf{X}^{(n)}}(q)-\sup_{q\in\mathcal{Q}_{K,\xi}}\triangle_{\mathbf{X}^{(n),\text{new}}}(q)\bigg{|}$
$\leq\sup_{q\in\mathcal{Q}_{K,\xi}}\bigg{|}\triangle_{\mathbf{X}^{(n)}}(q)-\triangle_{\mathbf{X}^{(n),\text{new}}}(q)\bigg{|}$
@@
Therefore, using McDiarmid’s inequality (10.5555/3153490 ), we have
@@
The covering number of $\mathcal{Q}_{K,\xi}$ can be easily derived using a similar approach used in Section C.4 of Supplementary Material:
$N(\epsilon,\mathcal{Q}_{K,\xi},\|\cdot\|_{\infty})$ $\leq n^{\phi_{j}}\bigg{(}1+{2\phi_{j}\xi\over\epsilon}\bigg{)}^{\phi_{j}}$
$\leq n^{K}\bigg{(}1+{2K\xi\over\epsilon}\bigg{)}^{K},$ (D.5)
where (D.5) is from $\phi_{j}\leq K$. Therefore, using Dudley Theorem (Theorem 1.19 in wolf2018MSbook ), we have
@@
Using Lemma D.3 in Section D.3 of Supplementary Material, we have
@@
$\geq 1-2^{|S|+1}|S|K^{|S|}\exp\bigg{(}-{K\log n\over 4\xi^{2}}\bigg{)}.$ (D.6)
The converse of Theorem D.1 is also true. That is, Theorem D.2 proves that the approximation error bettween a given populationally identifiable multinary-product tree and its empirically identifiable version is the same as that of Theorem D.1. The proof can be done by simply interchanging $\mu_{n,S}$ and $\mathbb{P}_{S}$ in the proof of Theorem D.1 and so is omitted.
Theorem D.2.
@@
$\mathbb{P}_{\mathbf{X}}^{n}\bigg{\{}\sup_{f_{S}\in\mathcal{F}^{P}_{S,K,\xi}}\|f_{S}-f_{\mu_{n,S},f_{S}}\|_{\infty}\leq(24\xi+1)2^{|S|}\sqrt{K\log n\over n}\bigg{\}}$
$\geq 1-2^{|S|+1}|S|K^{|S|}\exp\bigg{(}-{K\log n\over 4\xi^{2}}\bigg{)}.$
D.3 Rademacher complexity bound
Lemma D.3.
@@ D.3 Rademacher complexity bound
Appendix E Proof of Theorem 5.2
@@
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}+|\sigma^{2}-\sigma_{0}^{2}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0,$ (E.1)
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty$. (STEP 2). From (E.1), for any $S\subseteq[p]$ we establish that
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,\mathbb{P}_{\mathbf{X}}}+|\sigma^{2}-\sigma_{0}^{2}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0,$ (E.2)
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty$. (STEP 3). We modify (E.2) for the empirical $l_{2}$ norm. That is, we show that
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,n}+|\sigma^{2}-\sigma^{2}_{0}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0,$ (E.3)
@@
$\pi_{\xi}\big{\{}\Theta_{\xi}\backslash\Theta_{\xi}^{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0$ (E.4)
@@
E.1 Proof of (E.1)
We rely on the following result (See Theorem 19.3 of gyorfi2006distribution for its proof).
Lemma E.1 (Theorem 19.3 of gyorfi2006distribution ).
@@
$|g(\boldsymbol{x})|\leq K_{1},\quad\mathbb{E}[g(\boldsymbol{X})^{2}]\leq K_{2}\mathbb{E}[g(\boldsymbol{X})].$ (E.5)
@@
$\frac{\sqrt{n}\kappa(1-\kappa)t}{96\sqrt{2}\max\left\{K_{1},2K_{2}\right\}}\geq\int_{\frac{\kappa(1-\kappa)t}{16\max\left\{K_{1},2K_{2}\right\}}}^{\sqrt{t}}\sqrt{\log N\left(u,\left\{g\in\mathcal{G}:\frac{1}{n}\sum_{i=1}^{n}g\left(\mathbf{x}_{i}\right)^{2}\leq 16t\right\},||\cdot||_{1,n}\right)}du.$ (E.6)
@@
First, since $\mathcal{F}_{\xi}^{n}$ depends on the data $\mathbf{X}^{(n)}$, we cannot directly apply Lemma E.1. To resolve this problem, as in (D.4), we consider a class of linear combinations of binary product trees not necessarily identifiable, which we denote $\mathcal{F}_{\xi,\text{general}}^{n},$ and show that the covering number is of the same order as that of $\mathcal{F}_{\xi}^{n}$.
For given $S\subseteq[p],\mathbf{s}=(s_{j},j\in S)$, and $\boldsymbol{\beta}=(\beta_{\mathbf{v}},\mathbf{v}\in\{-1,1\}^{|S|})$, let $\mathbb{T}^{G}(\mathbf{x}_{S}:S,\mathbf{s},\boldsymbol{\beta})$ be a binary-product tree given as
$\mathbb{T}^{G}(\mathbf{x}_{S}:S,\mathbf{s},\boldsymbol{\beta})=\sum_{\mathbf{v}\in\{-1,1\}^{|S|}}\beta_{\mathbf{v}}\prod_{j\in S}\mathbb{I}^{(v_{j})}(x_{j}-s_{j}>0).$
@@
$\mathcal{F}_{\text{general}}^{n}=\Bigg{\{}f:f(\mathbf{x})=\sum_{t=1}^{T}\mathbb{T}^{G}\big{(}\mathbf{x}_{S_{t}}:S_{t},\mathbf{s}_{t},\boldsymbol{\beta}_{t}\big{)},\quad T\in[M_{n}]\>\>\>\text{and}\>\>\max_{t}\|\boldsymbol{\beta}_{t}\|_{\infty}\leq n^{p+1}\Bigg{\}},$
and $\mathcal{F}_{\xi,\text{general}}^{n}=\{f\in\mathcal{F}_{\text{general}}^{n},\|f\|_{\infty}\leq\xi\}$. Note that $\mathcal{F}^{n}_{\xi}\subseteq\mathcal{F}^{n}_{\xi,\text{general}}$. By a similar approach used in to Section C.4 of Supplementary Material, we can show
$N(\epsilon,\mathcal{F}_{\xi,\text{general}}^{n},\|\cdot\|_{\infty})\lesssim M_{n}n^{M_{n}p}\bigg{(}1+{n^{p+1}M_{n}2^{p+1}\over\epsilon}\bigg{)}^{M_{n}2^{p}}.$ (E.7)
For $K_{1}=K_{2}=4\xi^{2}$, $\kappa:={1\over 2}$, $\zeta=\epsilon_{n}^{2}$, and $\mathcal{G}=\{g:g=(f_{0}-f)^{2},f\in\mathcal{F}_{\xi,\text{general}}^{n}\}$, we first verify Condition (E.5) and Condition (E.6) in Lemma E.1. Verifying Condition (E.5): Since $\|f\|_{\infty}\leq\xi$ for $f\in\mathcal{F}_{\xi,\text{general}}^{n}$, we have
@@
for all $g\in\mathcal{G}.$ Verifying Condition (E.6): Since
@@
$\lesssim{n\epsilon_{n}^{2}\over\log n}n^{n\epsilon_{n}^{2}p/\log n}\bigg{(}1+{4\xi n^{p+1}\epsilon_{n}^{2}2^{p+1}\over u\log n}\bigg{)}^{n\epsilon_{n}^{2}/\log n}$ (E.8)
for any $u>0$, where (E.8) is derived from (E.7). Therefore, for $t\geq{\epsilon_{n}^{2}/8},\>{t/4\over 16\max(K_{1},2K_{2})}\leq u\leq\sqrt{t}$, we have
$\log N(u,\mathcal{G},\|\cdot\|_{1,n})$ $\lesssim\log{n\epsilon_{n}^{2}\over\log n}+{n\epsilon_{n}^{2}\over\log n}p\log n+{n\epsilon_{n}^{2}\over\log n}\log\bigg{(}1+{4\xi n^{p+1}\epsilon_{n}^{2}2^{p+1}\over u\log n}\bigg{)}$
@@
Proof of Condition (E.1): From Lemma E.1, we obtain the following bound:
@@
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}+|\sigma^{2}-\sigma^{2}_{0}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}$
$\leq\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f-f_{0}\|_{2,n}+|\sigma^{2}-\sigma^{2}_{0}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}$
@@
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}+|\sigma^{2}-\sigma^{2}_{0}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0$
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty,$ which completes the proof of (E.1). ∎
E.2 Proof of (E.2)
@@
$f(\mathbf{x})=\sum_{S\subseteq[p]}f_{S}(\mathbf{x}_{S}),$
where $f_{S}$ is the ensemble of $T_{S}$ many identifiable binary-product trees. Let $K_{S}=\prod_{j\in S}|\mathcal{P}_{j}|$, where $\mathcal{P}_{j}s$ are partitions in $f_{S}$. Since $T_{S}\leq\sum_{S\subseteq[p]}T_{S}\leq M_{n}$, it follows that
@@
Therefore by Theorem D.1, we have
@@
For notational simplicity, we denote $f_{\mathbb{P}_{S},f_{S}}$ by $f_{S}^{P}$. Let $f^{P}=\sum_{S\subseteq[p]}f_{S}^{P}$. For $f^{P}$, we have
$\|f^{P}-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}^{2}$ $=\int_{\mathcal{X}}\bigg{\{}\sum_{S\subseteq[p]}(f_{S}^{P}(\mathbf{x}_{S})-f_{0,S}(\mathbf{x}_{S}))\bigg{\}}^{2}\mathbb{P}_{\mathbf{X}}(d\mathbf{x})$
$\geq{1\over C_{L}}\int_{\mathcal{X}}\bigg{\{}\sum_{S\subseteq[p]}(f_{S}^{P}(\mathbf{x}_{S})-f_{0,S}(\mathbf{x}_{S}))\bigg{\}}^{2}\prod_{j=1}^{p}\mathbb{P}_{j}(dx_{j})$
$={1\over C_{L}}\sum_{S\subseteq[p]}\int_{\mathcal{X}}(f_{S}^{P}(\mathbf{x}_{S})-f_{0,S}(\mathbf{x}_{S}))^{2}\prod_{j=1}^{p}\mathbb{P}_{j}(dx_{j})$ (E.9)
$\geq{C_{U}\over C_{L}}\sum_{S\subseteq[p]}\|f_{S}^{P}-f_{0,S}\|_{2,\mathbb{P}_{\mathbf{X}}}^{2}$
@@
for all $S\subseteq[p]$, where the equality (E.9) holds since $f_{S}^{P}$s satisfy the populational identifiability condition. Thus, we obtain the following lower bound for $\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}$ :
@@
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,\mathbb{P}_{\mathbf{X}}}+|\sigma^{2}-\sigma^{2}_{0}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0,$
@@
E.3 Proof of (E.3)
In the same manner as in (E.1), using Lemma E.1 with $\mathcal{G}=\{g:g=(f_{0,S}-f_{S})^{2},f\in\mathcal{F}_{\xi,\text{general}}^{n}\}$, we can obtain the following :
$\pi_{\xi}\big{\{}(f,\sigma^{2})\in\Theta_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,n}+|\sigma^{2}-\sigma^{2}_{0}|>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0$

### v3

@@ C.1 Identifiable transformation of multinary-product tree
$\mathbb{E}_{\nu,W}(\gamma_{\boldsymbol{\ell}}):=\sum_{\boldsymbol{\ell}_{W}\in\prod_{t\in W}[\phi_{t}]}\gamma_{(\boldsymbol{\ell}_{W^{c}},\boldsymbol{\ell}_{W})}\nu_{W}^{\text{ind}}\{I_{\boldsymbol{\ell}_{W}}\}.$ (C.1)
@@
$f_{\nu,f_{S}}(\mathbf{x}_{S})=\sum_{\boldsymbol{\ell}}\tilde{\gamma}_{\boldsymbol{\ell}}\mathbb{I}(\mathbf{x}_{S}\in I_{\boldsymbol{\ell}}).$ (C.2)
C.2 Approximation error between the empirically and populationally identifiable multinary-product trees
We first investigate the approximation error between a given empirically identifiable multinary-product tree and its populationally identifiable version obtained by the formula (C.2), whose result is given in the following theorem. For $S\in\mathbb{S}$, let $\mathbb{P}_{S}$ be the distribution of $\mathbf{X}_{S}$ and $\mathbb{P}_{S}^{\text{ind}}=\prod_{j\in S}\mathbb{P}_{j}$, where $\mathbb{P}_{j}$ is the probability distribution of $\mathbf{X}_{j}$.
Theorem C.1.
@@ C.2 Approximation error between the empirically and populationally identifiable multinary-product trees
$=\sum_{\ell_{j}\in[\phi_{j}]}\gamma_{(\boldsymbol{\ell}_{j^{c}},\ell_{j})}\bigg(\mathbb{P}_{j}\{I_{j,\ell_{j}}\}-{1\over n}\sum_{i=1}^{n}\mathbb{I}(X_{i,j}\in I_{j,\ell_{j}})\bigg).$ (C.3)
In the following, we will show that (C.3) is small.
@@
$\mathbb{P}_{\mathbf{X}}^{n}\bigg\{\sup_{q\in\mathcal{Q}_{K,\xi}}\bigg|{1\over n}\sum_{i=1}^{n}q(X_{i,j})-\mathbb{E}_{j}[q(X_{j})]\bigg|>(24\xi+1)\sqrt{K\log n\over n}\bigg\}\leq 2\exp\bigg(-{K\log n\over 4\xi^{2}}\bigg),$ (C.4)
where $\mathcal{Q}_{K,\xi}$ is the set of all $q_{\mathcal{P},\boldsymbol{\gamma}}$ with $\phi_{j}\leq K$ and $\boldsymbol{\gamma}\in[-\xi,\xi]^{\phi_{j}},$ and $\mathbb{E}_{j}$ denotes the expectation under $\mathbb{P}_{j}$. Here, the interval partitions $I_{\ell}\subseteq[0,1]$ for $q\in\mathcal{Q}_{K,\xi}$ are arbitrary subsets of $[0,1]$ that do not depend on $\mathbf{X}^{(n)}$.
@@
$\bigg|\sup_{q\in\mathcal{Q}_{K,\xi}}\triangle_{\mathbf{X}^{(n)}}(q)-\sup_{q\in\mathcal{Q}_{K,\xi}}\triangle_{\mathbf{X}^{(n),\text{new}}}(q)\bigg|$ $\leq\sup_{q\in\mathcal{Q}_{K,\xi}}\bigg|\triangle_{\mathbf{X}^{(n)}}(q)-\triangle_{\mathbf{X}^{(n),\text{new}}}(q)\bigg|$
@@
Therefore, using McDiarmid’s inequality (Sammut and Webb (2017)), we have
@@
Let $\mathcal{Q}_{K,\xi,\text{order}}$ denote the function class consisting of functions $q(\cdot)$ defined above, whose split values are determined using the order statistics of $\mathbf{X}^{(n)}$ as in (7). Since $N(\epsilon,\mathcal{Q}_{K,\xi},\|\cdot\|_{2,n})=N(\epsilon,\mathcal{Q}_{K,\xi,\text{order}},\|\cdot\|_{2,n})$, the covering number of $\mathcal{Q}_{K,\xi}$ is easily derived using a similar approach used in Section B.4 of Supplementary Material:
$N(\epsilon,\mathcal{Q}_{K,\xi},\|\cdot\|_{2,n})$ $\leq n^{\phi_{j}}\bigg(1+{2\phi_{j}\xi\over\epsilon}\bigg)^{\phi_{j}}$
$\leq n^{K}\bigg(1+{2K\xi\over\epsilon}\bigg)^{K},$ (C.5)
where (C.5) is from $\phi_{j}\leq K$. Therefore, using Dudley Theorem (Theorem 1.19 in Wolf (2018)), we have
@@
Using Lemma C.3 in Section C.3 of Supplementary Material, we have
@@
$\geq 1-2^{|S|+1}|S|K^{|S|}\exp\bigg(-{K\log n\over 4\xi^{2}}\bigg).$ (C.6)
The converse of Theorem C.1 is also true. That is, Theorem C.2 proves that the approximation error bettween a given populationally identifiable multinary-product tree and its empirically identifiable version is the same as that of Theorem C.1. The proof can be done by simply interchanging $\mu_{n,S}$ and $\mathbb{P}_{S}$ in the proof of Theorem C.1 and so is omitted.
Theorem C.2.
@@
$\mathbb{P}_{\mathbf{X}}^{n}\bigg\{\sup_{f_{S}\in\mathcal{F}^{P}_{S,K,\xi}}\|f_{S}-f_{\mu_{n,S},f_{S}}\|_{\infty}\leq(24\xi+1)2^{|S|}\sqrt{K\log n\over n}\bigg\}\geq 1-2^{|S|+1}|S|K^{|S|}\exp\bigg(-{K\log n\over 4\xi^{2}}\bigg).$
C.3 Rademacher complexity bound
Lemma C.3.
@@ C.3 Rademacher complexity bound
Appendix D Proof of Theorem 5.2
@@
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0,$ (D.1)
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty$. (STEP 2). From (D.1), for any $S\in\mathbb{S}$ we establish that
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,\mathbb{P}_{\mathbf{X}}}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0,$ (D.2)
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty$. (STEP 3). We modify (D.2) for the empirical $l_{2}$ norm. That is, we show that
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,n}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0,$ (D.3)
@@
$\pi_{\xi}\big\{\mathcal{F}_{\xi}\backslash\mathcal{F}_{\xi}^{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0$ (D.4)
@@
D.1 Proof of (D.1)
We rely on the following result (See Theorem 19.3 of Györfi et al. (2006) for its proof).
Lemma D.1 (Theorem 19.3 of Györfi et al. (2006)).
@@
$|g(\boldsymbol{x})|\leq K_{1},\quad\mathbb{E}[g(\boldsymbol{X})^{2}]\leq K_{2}\mathbb{E}[g(\boldsymbol{X})].$ (D.5)
@@
$\frac{\sqrt{n}\kappa(1-\kappa)t}{96\sqrt{2}\max\left\{K_{1},2K_{2}\right\}}\geq\int_{\frac{\kappa(1-\kappa)t}{16\max\left\{K_{1},2K_{2}\right\}}}^{\sqrt{t}}\sqrt{\log N\left(u,\left\{g\in\mathcal{G}:\frac{1}{n}\sum_{i=1}^{n}g\left(\mathbf{x}_{i}\right)^{2}\leq 16t\right\},||\cdot||_{1,n}\right)}du.$ (D.6)
@@
First, since $\mathcal{F}_{\xi}^{n}$ depends on the data $\mathbf{X}^{(n)}$, we cannot directly apply Lemma D.1. To resolve this problem, as in (C.4), we consider a class of linear combinations of binary product trees not necessarily identifiable, with split values chosen as scalars in $[0,1]$ that do not depend on $\mathbf{X}^{(n)}$. We denote this class by $\mathcal{F}_{\xi,\text{general}}^{n}$ and show that its covering number is of the same order as that of $\mathcal{F}_{\xi}^{n}$.
For given $S\subseteq[p],\>\mathbf{s}=(s_{j}\in[0,1],j\in S)$, and $\boldsymbol{\beta}=(\beta_{1,j},\beta_{2,j},j\in S)$, let $\mathbb{T}^{G}(\mathbf{x}_{S}:S,\mathbf{s},\boldsymbol{\beta})$ be a binary-product tree given as
$\mathbb{T}^{G}(\mathbf{x}_{S}:S,\mathbf{s},\boldsymbol{\beta})=\prod_{j\in S}\bigg(\beta_{1,j}\mathbb{I}(x_{j}-s_{j}\leq 0)+\beta_{2,j}\mathbb{I}(x_{j}-s_{j}>0)\bigg).$
@@
$\mathcal{F}_{\text{general}}^{n}=\Bigg\{f:f(\mathbf{x})=\sum_{t=1}^{T}\mathbb{T}^{G}\big(\mathbf{x}_{S_{t}}:S_{t},\mathbf{s}_{t},\boldsymbol{\beta}_{t}\big),\quad T\in[M_{n}],\>\mathbf{s}_{t}\in[0,1]^{|S_{t}|},\max_{t}\|\boldsymbol{\beta}_{t}\|_{\infty}\leq n\Bigg\},$
and $\mathcal{F}_{\xi,\text{general}}^{n}=\{f\in\mathcal{F}_{\text{general}}^{n},\|f\|_{\infty}\leq\xi\}$. Note that $\mathcal{F}^{n}_{\xi}\subseteq\mathcal{F}^{n}_{\xi,\text{general}}$. By a similar approach used in Section B.4 of Supplementary Material and (C.5), we can show
$N(\epsilon,\mathcal{F}_{\xi,\text{general}}^{n},\|\cdot\|_{1,n})\lesssim M_{n}n^{M_{n}p}\bigg(1+{8np(n+1)^{p}M_{n}\over\epsilon}\bigg)^{4pM_{n}}.$ (D.7)
For $K_{1}=K_{2}=4\xi^{2}$, $\kappa:={1\over 2}$, $\zeta=\epsilon_{n}^{2}$, and $\mathcal{G}=\{g:g=(f_{0}-f)^{2},f\in\mathcal{F}_{\xi,\text{general}}^{n}\}$, we first verify Condition (D.5) and Condition (D.6) in Lemma D.1. Verifying Condition (D.5): Since $\|f\|_{\infty}\leq\xi$ for $f\in\mathcal{F}_{\xi,\text{general}}^{n}$, we have
@@
for all $g\in\mathcal{G}.$ Verifying Condition (D.6): Since
@@
$\lesssim{n\epsilon_{n}^{2}\over\log n}n^{n\epsilon_{n}^{2}p/\log n}\bigg(1+{32\xi np(1+n)^{p}\epsilon_{n}^{2}\over u\log n}\bigg)^{n\epsilon_{n}^{2}/\log n}$ (D.8)
for any $u>0$, where (D.8) is derived from (D.7). Therefore, for $t\geq{\epsilon_{n}^{2}/8},\>{t/4\over 16\max(K_{1},2K_{2})}\leq u\leq\sqrt{t}$, we have
$\log N(u,\mathcal{G},\|\cdot\|_{1,n})$ $\lesssim\log{n\epsilon_{n}^{2}\over\log n}+{n\epsilon_{n}^{2}\over\log n}p\log n+{n\epsilon_{n}^{2}\over\log n}\log\bigg(1+{32\xi np(1+n)^{p}\epsilon_{n}^{2}\over u\log n}\bigg)$
@@
Proof of Condition (D.1): From Lemma D.1, we obtain the following bound:
@@
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}$
$\leq\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f-f_{0}\|_{2,n}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}$
@@
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0$
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty,$ which completes the proof of (D.1). ∎
D.2 Proof of (D.2)
@@
$f(\mathbf{x})=\sum_{S\in\mathbb{S}}f_{S}(\mathbf{x}_{S}),$
where $f_{S}$ is the ensemble of $T_{S}$ many identifiable binary-product trees. Let $K_{S}=\prod_{j\in S}|\mathcal{P}_{j}|$, where $\mathcal{P}_{j}s$ are partitions in $f_{S}$. Since $T_{S}\leq\sum_{S\in\mathbb{S}}T_{S}\leq M_{n}$, it follows that
@@
Therefore by Theorem C.1, we have
@@
For notational simplicity, we denote $f_{\mathbb{P}_{S},f_{S}}$ by $f_{S}^{P}$. Let $f^{P}=\sum_{S\in\mathbb{S}}f_{S}^{P}$. For $f^{P}$, we have
$\|f^{P}-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}^{2}$ $=\int_{\mathcal{X}}\bigg\{\sum_{S\in\mathbb{S}}(f_{S}^{P}(\mathbf{x}_{S})-f_{0,S}(\mathbf{x}_{S}))\bigg\}^{2}\mathbb{P}_{\mathbf{X}}(d\mathbf{x})$
$\geq{1\over C_{L}}\int_{\mathcal{X}}\bigg\{\sum_{S\in\mathbb{S}}(f_{S}^{P}(\mathbf{x}_{S})-f_{0,S}(\mathbf{x}_{S}))\bigg\}^{2}\prod_{j=1}^{p}\mathbb{P}_{j}(dx_{j})$
$={1\over C_{L}}\sum_{S\in\mathbb{S}}\int_{\mathcal{X}}(f_{S}^{P}(\mathbf{x}_{S})-f_{0,S}(\mathbf{x}_{S}))^{2}\prod_{j=1}^{p}\mathbb{P}_{j}(dx_{j})$ (D.9)
$\geq{C_{U}\over C_{L}}\sum_{S\in\mathbb{S}}\|f_{S}^{P}-f_{0,S}\|_{2,\mathbb{P}_{\mathbf{X}}}^{2}$
@@
for all $S\in\mathbb{S}$, where the equality (D.9) holds since $f_{S}^{P}$s satisfy the populational identifiability condition. Thus, we obtain the following lower bound for $\|f-f_{0}\|_{2,\mathbb{P}_{\mathbf{X}}}$ :
@@
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,\mathbb{P}_{\mathbf{X}}}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0,$
@@
D.3 Proof of (D.3)
In the same manner as in (D.1), using Lemma D.1 with $\mathcal{G}=\{g:g=(f_{0,S}-f_{S})^{2},f\in\mathcal{F}_{\xi,\text{general}}^{n}\}$, we can obtain the following :
$\pi_{\xi}\big\{f\in\mathcal{F}_{\xi}^{n}:\|f_{S}-f_{0,S}\|_{2,n}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0$